OpenStudy (anonymous):

what is the asymptote for the hyperbola... ? (y+1)^2 (x+4)^2 ________ _ ________ 300 48

OpenStudy (anonymous):

can anyone e help?

OpenStudy (amistre64):

\[y-C_y=\pm\sqrt{\frac{b^2}{a^2}}(x-C_x)\]

OpenStudy (anonymous):

so... would it be \[\pm(1/4)(x+4)+1\] ?

OpenStudy (amistre64):

since this thing never touches the "x" axis; a^2 is under the y part and the b^2 is under the x part;

OpenStudy (amistre64):

is that a guess from the answer options? or did you try to work that out?

OpenStudy (anonymous):

i tried to figure it out... but i really don't understand...

OpenStudy (amistre64):

the asymps run thru the center so we want that point on our line; the slope is defined as .... hmm, sqrt(ypart/xpart) now that i think of this

OpenStudy (amistre64):